How to find the angle{theta} when the question is given like this


que: two equal discs initially at rest are in contact on a table. a third disc of same mass but of double radius strikes them symmetrically and itself comes to rest after impact.the coeficient of restitution is


1] 2/3


2] 9/16


3] 3/4


4] 8/17

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

let speed before collision be u, and after collision be v. conservation of linear momentum in the direction of motion of the bigger disk gives 2mv(2√2)/3=mu, v=3u/4√2, let 'theta' be the angle between u and v. coefficient of restitution

=v/(u(cosθ))=(3u/4√2)/(u(√8)/3)=9/16

'Win exciting gifts by answering the questions on Discussion Forum.'